### v1

# Realistic Extreme Image Rescaling via Generative Latent Space Learning

Abstract
Image rescaling aims to learn the optimal downscaled low-resolution (LR) image that can be accurately reconstructed to its original high-resolution (HR) counterpart. This process is crucial for efficient image processing and storage, especially in the era of ultra-high definition media. However, extreme downscaling factors pose significant challenges due to the highly ill-posed nature of the inverse upscaling process, causing existing methods to struggle in generating semantically plausible structures and perceptually rich textures. In this work, we propose a novel framework called Latent Space Based Image Rescaling (LSBIR) for extreme image rescaling tasks. LSBIR effectively leverages powerful natural image priors learned by a pre-trained text-to-image diffusion model to generate realistic HR images. The rescaling is performed in the latent space of a pre-trained image encoder and decoder, which offers better perceptual reconstruction quality due to its stronger sparsity and richer semantics. LSBIR adopts a two-stage training strategy. In the first stage, a pseudo-invertible encoder-decoder models the bidirectional mapping between the latent features of the HR image and the target-sized LR image. In the second stage, the reconstructed features from the first stage are refined by a pre-trained diffusion model to generate more faithful and visually pleasing details. Extensive experiments demonstrate the superiority of LSBIR over previous methods in both quantitative and qualitative evaluations. The code will be available at: https://github.com/wwangcece/LSBIR.
1 Introduction
With the explosive growth of ultra high-resolution digital visual data, image rescaling has become essential for enhancing data storage and transmission efficiency, optimizing resource usage in real-time applications, and ensuring proper display on devices with varying resolutions. During image rescaling, high-resolution (HR) images are often downscaled to visually pleasing low-resolution (LR) images for transmission or display. Subsequently, an upscaling process is employed to restore the downscaled images to their original resolution. Unlike the super-resolution (SR) task, image rescaling allows for joint optimization of both downscaling and upscaling processes, preserving detailed information crucial for effective upscaling.
Figure 1: Visual comparison of 16$\times$ image rescaling results. HCFlow struggles to obtain sufficient texture, while the reconstruction results of VQIR lack accurate semantic and structural information.
In recent years, various image rescaling methods have been proposed, which can be classified into two main architectures: encoder-decoder [11, 22] and invertible neural networks [30, 16]. The encoder-decoder architecture employs two sets of neural networks to perform downscaling and upscaling operations. In contrast, invertible neural networks view image rescaling as a reversible process, using the forward and backward passes of an invertible neural network to simulate the downscaling and upscaling operations. Despite the significant improvements in reconstruction quality achieved by these methods, they are limited to small rescaling factors, such as 2$\times$ or 4$\times$.
However, with the advancement of HR imaging devices, there is an increasing need for extreme rescaling factors (e.g., 16$\times$ or 32$\times$) to achieve more efficient storage and transmission of ultra HR visual data. When dealing with extreme rescaling factors, existing methods often struggle to generate sufficient details. Therefore, some approaches propose using pre-trained GANs as priors to constrain the solution space of the reconstructed results. GRAIN [36] and BDFlow [13] utilize pre-trained StyleGAN [8, 9] to generate high-quality upscaling face images. VQIR [29], on the other hand, leverages the high-quality visual embeddings encapsulated in pre-trained VQGAN [4] to achieve extreme rescaling on natural images. Although these methods have made significant progress in perceptual metrics, we find that their reconstruction results suffer from shortcomings in overall semantics and structure. As shown in Fig. 1, while VQIR generates richer details compared to HCFlow, its reconstruction results lack precise semantic and structural information. This is due to the use of both latent space and pixel space losses. The introduction of pixel-space loss can force the model to compromise on both optimization paths, leading to the generation of significant artifacts.
Figure 2: Image rescaling within pixel-space and latent-space. Different from methods that impose pixel loss, our approach applies loss only in the latent space, thereby ensuring the semantic integrity of the reconstruction results.
Therefore, we propose a novel framework named Latent Space Based Image Rescaling (LSBIR) for extreme rescaling, which does not impose any pixel loss on the reconstruction results, as illustrated in Fig. 2. Rescaling operations in the latent space have a broader receptive field, allowing richer information to be embedded into the LR images. Additionally, training in the latent space preserves the semantic integrity of the resampled features [7], better aligning with the intrinsic priors of pre-trained generative models. Specifically, LSBIR adopts a two-stage training strategy. In the first stage, a simple convolutional neural network (CNN) is used to rescale the latent code of the HR images. Concurrently, an invertible neural network (INN) completes the bidirectional mapping between the feature domain and the pixel domain. In the second stage, the rescaled features from the first stage guide the generation process of the pre-trained Stable Diffusion (SD) model, enhancing the details of the rescaled features. The first stage ensures that the rescaled features retain complete semantic information of the HR images, while the second stage generates richer textures and improves perceptual quality, balancing fidelity and realism in the final results.
Our contributions can be summarized as follows:
  • We propose a Latent Space Based Image Rescaling framework for extreme rescaling. Without imposing any pixel loss on the reconstruction results, our method exhibits excellent performance in global semantic and structural reconstruction.
  • We propose a novel decoupled strategy for feature rescaling. By decoupling the feature-pixel transform operation from the rescaling operation, this module can enhance the fidelity of feature reconstruction.
  • To the best of our knowledge, we are the first to apply diffusion models to the field of extreme image rescaling. We demonstrate that the rich knowledge contained in the pre-trained generative diffusion model can significantly enhance the perceptual quality of rescaling results.
Figure 3: Overview of proposed Latent-Space Based Image Rescaling (LSBIR). In the first training stage, we train a feature rescaling network in the latent space of a pre-trained VAE, which outputs LR images and rescaled features. In the second training stage, we use the features from the first stage as conditions and train an adapter to guide the generation process of the pre-trained SD model.
2 Related work
2.1 Image Rescaling
Image rescaling aims to downscale HR images to visually coherent LR images and then plausibly reconstruct the original HR images. Unlike image SR, image rescaling involves joint modeling of both downscaling and upscaling processes, leading to higher reconstruction accuracy. Earlier works, such as those by [11, 15], utilize CNNs to build an encoder-decoder framework, treating image rescaling as a unified task by jointly training the downscaling and upscaling processes. Instead of directly generating LR images, CAR [22] proposes generating content-adaptive resampling kernels based on HR images to perform the downsampling operation, thus avoiding pixel-level constraints on the downscaled images. To fully leverage the inherent reversibility of image rescaling, there has been a surge in the use of INNs for this task in recent years [14, 37, 33, 31]. IRN [30] is the first invertible framework that models image rescaling as a bijective transformation, embedding residual high-frequency components into a case-agnostic latent distribution for efficient reconstruction. HCFlow [16] assumes that high-frequency (HF) information depends on the LR image, thereby achieving better performance by incorporating LR conditions in HF information estimation. In the realm of extreme image rescaling, generative priors are usually employed to constrain the outcomes. Notably, GRAIN [36] and VQIR [29] leverage pre-trained StyleGAN [8, 9] and VQGAN [4], respectively, to achieve notable enhancements in perceptual quality. Nonetheless, these approaches are often limited in scope, either confined to handling images from specific domains or encountering difficulties when reconstructing faces and textual elements.
Recently, the emergence of the pre-trained SD model [20] has showcased its efficacy in both content generation [32] and image restoration [25]. Despite this progress, to date, no effort has been made to integrate it within the framework of extreme image rescaling, presenting an unexplored avenue with potential for advancing the field.
2.2 Diffusion Model for Image SR
Here, we briefly introduce the application of diffusion models in a related task, i.e., image SR. Diffusion models promote the development of natural image SR in two main ways: training diffusion models from scratch and leveraging pre-trained SD model. The former is represented by SR3 [21], which achieves perceptual quality comparable to GANs. The latter is represented by StableSR [25], which aims to build a model capable of handling image restoration under any complex degradation in real-world scenarios.
However, due to the inherent difficulty of integrating diffusion models into existing encoder-decoder architectures, few work has yet attempted to combine them with extreme image rescaling. In this work, we decompose extremec image rescaling into two sub-tasks: feature rescaling in the latent space and perception-oriented optimization. By doing so, we could explore the potential application of diffusion models in this field.
3 Methodology
3.1 Overview of LSBIR
Our LSBIR architecture, shown in Fig. 3, involves two training stages: Feature Rescaling in the Latent Space and Perception-oriented Optimization. Initially, the input HR image is encoded into a latent feature $z$, which is then rescaled by the Decoupled Feature Rescaling Module (DFRM) to generate an LR image $y$ and a rescaled latent feature $\hat{z}$. The first stage aims to align $\hat{z}$ with $z$ while ensuring $y$ retains the content of the original HR image, preserving semantic information for the second stageś refinement. Although decoding $\hat{z}$ produces high-fidelity HR images, they lack texture detail. Thus, in the second stage, we use an adapter to guide the pre-trained SD model, generating more realistic textures. During inference, we could balance fidelity and realism by weighting $\hat{z}$ from both stages.
3.2 Feature Rescaling in the Latent Space
3.2.1 Overall Network Architecture
In the first stage, we perform rescaling of the HR features $z$ in the latent space, obtaining LR images $y$ and rescaled features $\hat{z}$. A straightforward idea is to construct a transformation chain as $y=G_{e}\left(z\right)$, $\hat{z}=G_{d}\left(y\right)$. Here, $G_{e}$ is an encoder used to downscale the input features to the target size, while $G_{d}$ represents the decoder used to reconstruct the original high-resolution features. Guidance loss $\mathcal{L}_{gui.}$ and reconstruction loss $\mathcal{L}_{rec.}$ are applied to $y$ and $\hat{z}$ respectively, to force $y$ to have the same content as $x$, and $\hat{z}$ to be as similar as possible to $z$:
$\mathcal{L}_{\mathrm{gui.}}=\parallel y-\mathrm{Bicubic}(x)\parallel_{p}$ (1)
$\mathcal{L}_{\mathrm{rec.}}=\parallel\hat{z}-z\parallel_{p}$ (2)
However, these two losses are contradictory to each other. When the guidance loss approaches zero, the rescaling problem degenerates into a classical SR [28] problem. Conversely, when the reconstruction loss reaches its global optimum, the rescaling problem degenerates into an image compression problem [18]. Therefore, directly applying the guidance loss $\mathcal{L}_{gui.}$ on $y$ might negatively impact subsequent upscaling.
Considering the above issues, we devise the DFRM as shown in Fig. 4. Inspired by FGRN [14], we use a set of CNNs to perform downscaling and upscaling, and a separate set of INNs to perform the bidirectional mapping between the feature domain and the pixel domain. To achieve balance between feature rescaling accuracy and visual quality of LR images, we design another pixel guidance module in the encoder to incorporate pixel-level information. Specifically, DFRM perform feature rescaling through two independent transformation chains $\left(\left(x,z\right)\rightarrow w\rightarrow\hat{z}\right)$; $\left(w\leftrightarrow y\right)$ as follows:
$w=G_{e}\left(x,z\right)$ (3)
$w\approx F^{-1}\left(F\left(w\right)\right)=F^{-1}\left(y\right)$ (4)
$\hat{z}=G_{d}\left(w\right)$ (5)
Here, $F$ and $F^{-1}$ represent the forward and reverse processes of the INN, respectively. Given the latent variable $z\in R^{H/8\times W/8\times 4}$ of the pre-trained SD VAE as input, the encoder $G_{e}$ first transforms it into a compact representation $w\in R^{H/N\times W/N\times 3}$, whose size is the same as that of the target LR image and $N$ represents the target downscaling factor. Subsequently, a set of INNs perform the forward process to transform $w$ from the feature domain to the pixel domain and obtain the LR image $y\in R^{H/N\times W/N\times 3}$ through quantization. Symmetrically, in the upscaling process, the inverse of the INN is used to process the LR image $y$ and reconstruct $w$. Finally, the decoder $G_{d}$ is utilized to obtain the rescaled feature $\hat{z}$.
Figure 4: Detailed architecture of Decoupled Feature Rescaling Module (DFRM).
3.2.2 Domain Converter
The domain converter module consists of a series of stacked invertible blocks [3], which can be defined as $F=f_{1}\circ f_{2}\circ...f_{K}$. For the $i$-th invertible block $f_{i}$ in the forward process, the input $w_{k}$ is split into $w_{k}^{1}$ and $w_{k}^{2}$ along the channel axis, and they undergo affine coupling layers for information fusion as follows:
$w_{k+1}^{1}=w_{k}^{1}+\phi\left(w_{k}^{2}\right)$ (6)
$w_{k+1}^{2}=w_{k}^{2}\otimes\exp\left(\rho\left(w_{k+1}^{1}\right)\right)+\eta\left(w_{k+1}^{1}\right)$ (7)
Here, $\phi\left(\cdot\right)$, $\rho\left(\cdot\right)$ and $\eta\left(\cdot\right)$ could be arbitrary functions and we adopt the widely used RRDB [26] for its good representation ability. Finally, the output of the last invertible module, i.e., $f_{K}$, is quantized to obtain the LR image $y$. Similarly, in the reverse process, the LR image $y$ will be used as an input to reconstruct $w$. For the $i$-th invertible block $f_{i}$, the input $w_{k+1}$ is split into $w_{k+1}^{1}$ and $w_{k+1}^{2}$ and undergo affine coupling layers as follows, thus obtaining the low-resolution representation $w$:
$w_{k}^{2}=\left(w_{k+1}^{2}-\eta\left(w_{k+1}^{1}\right)\right)\otimes\exp\left(-\rho\left(w_{k+1}^{1}\right)\right)$ (8)
$w_{k}^{1}=w_{k+1}^{1}-\phi\left(w_{k}^{2}\right)$ (9)
3.2.3 Loss Function
First stage training involves two losses: reconstruction loss and guidance loss. Considering the information lost in quantization operations, the bidirectional mapping between $w$ and $y$ is not exactly reversible. To ensure the robustness of the network to quantization operations, we consider two types of transformation chains in designing the reconstruction loss: $\left(x,z\right)\rightarrow w\rightarrow\hat{z}$ and $\left(x,z\right)\rightarrow w\rightarrow y\rightarrow\hat{w}\rightarrow\hat{z}$, which can be expressed as the following formula:
$\begin{split}\mathcal{L}_{\mathrm{rec.}}=&\parallel G_{d}\left(G_{e}\left(x,z\right)\right)-z\parallel_{p}+\\ &\parallel G_{d}\left(F^{-1}\left(F\left(G_{e}\left(x,z\right)\right)\right)\right)-z\parallel_{p}\end{split}$ (10)
This combination can facilitate the network to reconstruct $\hat{z}$ with high-fidelity under both the existence and absence of quantization operations. Additionally, considering that the guidance loss should be decoupled from the reconstruction loss as much as possible, we design it as follows:
$\mathcal{L}_{\mathrm{gui.}}=\parallel\mathrm{sg}\left(F\left(G_{e}\left(x,z\right)\right)\right)-\mathrm{Bicubic}(x)\parallel_{p}\$ (11)
Here, $\mathrm{sg}\left(\cdot\right)$ is the stop-gradient operation preventing the gradient flowing into $G_{e}$ or $G_{d}$, which decouples the feature-pixel domain bidirectional mapping and feature rescaling operations. Finally, the loss in the first stage is a combination of these two losses:
$\mathcal{L}_{\mathrm{stage1}}=\lambda_{1}\mathcal{L}_{\mathrm{rec.}}+\lambda_{2}\mathcal{L}_{\mathrm{gui.}}\$ (12)
Table 1: Quantitative comparisons with different methods at 16$\times$ and 32$\times$. The symbols $\uparrow$ and $\downarrow$ respectively represent that higher or lower values indicate better performance. Bold represents the best and underline represents the second best.
Dataset Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ MUSIQ $\uparrow$ CLIPIQA $\uparrow$
16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$ 16$\times$ 32$\times$
DIV2K EDSR 24.64 22.13 0.6572 0.6056 0.5376 0.6219 0.3162 0.4184 34.15 23.28 0.3397 0.4193
RCAN 24.34 22.06 0.6479 0.6034 0.5485 0.6307 0.3269 0.4349 31.57 23.18 0.3408 0.4091
ESRGAN 23.15 21.15 0.5946 0.5808 0.4478 0.5788 0.2378 0.4237 59.80 43.40 0.6161 0.3725
IRN 26.04 22.84 0.7020 0.6212 0.5153 0.6095 0.3041 0.4036 43.72 28.70 0.2799 0.2336
HCFlow 26.66 23.89 0.7176 0.6467 0.4885 0.5816 0.2866 0.3852 46.43 37.25 0.2735 0.2792
VQIR 23.91 22.02 0.6498 0.5823 0.3174 0.4568 0.1024 0.2663 64.04 58.21 0.6350 0.6293
Ours 22.58 20.41 0.6171 0.5442 0.3559 0.4763 0.1188 0.1964 67.73 65.27 0.6934 0.6528
Urban100 EDSR 19.93 18.08 0.5020 0.4447 0.5827 0.6839 0.3701 0.4734 41.97 24.80 0.3439 0.4173
RCAN 19.96 18.14 0.5017 0.4452 0.5859 0.6880 0.3745 0.4808 40.31 24.46 0.3561 0.4130
ESRGAN 19.36 17.68 0.4835 0.4344 0.4741 0.6142 0.2554 0.4288 65.82 57.02 0.5849 0.3860
IRN 22.15 18.93 0.6155 0.4772 0.5114 0.6567 0.3338 0.4457 58.73 32.03 0.3217 0.2210
HCFlow 22.59 19.86 0.6335 0.5185 0.4841 0.6048 0.3125 0.4101 60.62 47.94 0.3182 0.2827
VQIR 20.27 18.44 0.5782 0.4427 0.3038 0.5045 0.1418 0.3290 71.56 61.64 0.6619 0.6568
Ours 22.16 17.76 0.6809 0.4342 0.2647 0.5164 0.1253 0.2904 72.71 69.72 0.6832 0.6602
CLIC2020 EDSR 26.33 23.92 0.7408 0.7062 0.5109 0.5742 0.3057 0.4024 33.17 21.77 0.3392 0.4307
RCAN 26.37 24.04 0.7407 0.7072 0.5133 0.5777 0.3126 0.4151 31.59 22.07 0.3469 0.4180
ESRGAN 25.43 23.48 0.7018 0.6991 0.4213 0.5347 0.2209 0.4014 57.80 40.98 0.6107 0.3537
IRN 28.55 24.93 0.7907 0.7228 0.4676 0.5528 0.2746 0.3776 47.66 27.81 0.2950 0.2495
HCFlow 28.82 26.21 0.7949 0.7451 0.4542 0.5325 0.2635 0.3611 47.72 37.15 0.2733 0.2884
VQIR 25.69 24.12 0.7363 0.6942 0.2854 0.4191 0.0901 0.2445 64.36 54.83 0.6105 0.6068
Ours 24.62 22.59 0.7063 0.6551 0.3423 0.4618 0.1217 0.2036 66.48 61.67 0.6662 0.6229
DIV8K EDSR 26.56 24.48 0.6776 0.6472 0.5437 0.6048 0.3358 0.4325 30.83 22.30 0.3701 0.4686
RCAN 26.63 24.57 0.6778 0.6481 0.5470 0.6085 0.3430 0.4449 29.93 22.52 0.3796 0.4502
ESRGAN 25.49 23.78 0.6247 0.6350 0.4628 0.5610 0.2518 0.4268 53.49 39.11 0.6401 0.4078
IRN 28.95 25.46 0.7402 0.6636 0.4932 0.5870 0.3044 0.4066 42.61 27.76 0.3216 0.2741
HCFlow 29.22 26.69 0.7470 0.6857 0.4793 0.5616 0.2932 0.3906 42.71 35.53 0.2948 0.3128
VQIR 25.79 24.26 0.6889 0.6350 0.3221 0.4457 0.1066 0.2628 56.76 51.44 0.5769 0.6203
Ours 24.39 22.94 0.6515 0.5943 0.3846 0.5058 0.1524 0.2451 58.28 55.68 0.6663 0.6231
3.3 Perception-oriented Optimization
The $\hat{z}$ in the first stage can be passed to the decoder $\mathcal{D}$ to obtain the rescaled HR image. Although this result has high fidelity, it lacks rich texture. Therefore, training in the second stage is designed to enhance perceptual quality. Inspired by relevant content generation [32, 19], we adopt the adapter architecture to guide the generation of the SD model. The adapter module takes the rescaled features $\hat{z}$ as input and extracts multi-scale features in a coarse-to-fine manner, allowing it to output multi-scale features of different sizes:
$\{F_{c}^{1},F_{c}^{2},...,F_{c}^{K}\}=\mathrm{Ada}(\hat{z})$ (13)
Subsequently, these features are element-wisely added to different layers of the denoising U-Net. Considering that image rescaling task places a higher emphasis on fidelity, we employ conditional control over both the encoding and decoding layers of the denoising U-Net:
$F_{\mathrm{enc}}^{i}=F_{\mathrm{enc}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (14)
$F_{\mathrm{dec}}^{i}=F_{\mathrm{dec}}^{i}+F_{c}^{i},\ \ i\in\{1,2,...,K\}$ (15)
Here, $F_{\mathrm{enc}}^{i}$ and $F_{\mathrm{dec}}^{i}$ represent the output features of the encoder or decoder of the denoising U-Net at the $i$-th layer. Finally, the SD model and adapter are optimized using the following loss function:
$\mathcal{L_{\mathrm{stage2}}}=\mathrm{\mathbb{E}}_{\left(z,\epsilon,\hat{z},t\right)}\parallel\epsilon-\epsilon_{\theta}\left(z_{t},t,\hat{z}\right)\parallel^{2}$ (16)
Figure 5: Qualitative comparisons of 16$\times$ (the 2nd and 3rd rows) and 32$\times$ (the 4th and 5th rows) rescaling results. Our method achieves higher semantic accuracy, such as more recognizable faces, clearer text, and more realistic structures.
Figure 6: Visual comparisons of 16$\times$ rescaling methods on the Urban100 dataset, including both HR images and corresponding edge maps. Our model is capable of restoring more regular structures and producing sharper, more accurate edges.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets and Evaluation Metrics
We train the proposed model using the DIV2K dataset [1] and the Flickr2K dataset [17]. Due to the low resolution of conventional benchmark datasets such as Set5, Set14, and BSD100, they are not suitable for extreme rescaling tasks [5]. Therefore, we utilized four high-resolution datasets: DIV2K, Urban100 [6], CLIC2020 [23], and DIV8K [5] to evaluate our model. We adopt PSNR and SSIM [27] for quantitative evaluation. For perceptual quality evaluation, reference-based metrics LPIPS [34] and DISTS [2] are used. For non-reference image quality assessment, we use metrics including MUSIQ [10] and CLIPIQA [24].
4.1.2 Training Details
Our LSDIR is built based on the SD 2.1-base model and the training is divided into two stages. First, we train the DFRM together with the pre-trained VAE model for 200K iterations using Eq. 12 The learning rate is initialized as $2\times 10^{-4}$ and reduced by half every 50k iterations. After the training of DFRM, we adopt Eq. (16) to train the Adapter and finetune the diffusion U-Net. We only finetune the linear layers of the U-Net for its better performance. During both stages, we adopt Adam optimizer [12] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and set the mini-batch size to 32.
4.2 Results and Comparison with SOTA
[…]
4.2.1 Quantitative Results
Tab. 1 presents the 16$\times$ and 32$\times$ quantitative image rescaling results on four datasets. The performance of SR methods are generally poor. While methods based on INNs achieve the highest PSNR and SSIM scores due to consideration of the correlation between downscaling and upscaling. VQIR exhibits the best LPIPS by fully utilizing the prior provided by pre-trained VQGAN and being trained based on LPIPS-loss. As our proposed method is trained in the latent space, the results exhibit the best semantic integrity, which leads to significant improvements in two non-reference metrics. It is important to note that while our method slightly lags behind VQIR in terms of certain full-reference metrics, it excels in reconstructing accurate semantics and regular structures, as will be demonstrated in the following section.
4.2.2 Qualitative Results
In Fig. 5, we present visual comparisons at 16$\times$ and 32$\times$ rescaling factors. Due to the ill-posed nature of extreme rescaling, pixel-oriented optimization methods such as EDSR, RCAN, IRN, and HCFlow produce overly-smoothed results. Although ESRGAN incorporates perceptual optimization it still cannot generate adequate details. VQIR leverages the pre-trained VQGAN as a prior, allowing it to generate rich details and sharp edges. However, it struggles with faces and more complex structures. Thanks to the powerful prior provided by the pre-trained SD and the rich semantic information contained in the latent space of the pre-trained VAE, our method effectively addresses these issues. It not only generates sufficiently rich textures but also reconstructs accuracy semantics and regular structures, such as faces, text and architectural structures.
The quantitative experiments demonstrate that our model exhibits outstanding performance on the Urban100 dataset, confirming that our method produces more accurate structural features. Therefore, we provide a visual comparison of 16$\times$ rescaling results on the Urban100 dataset, as shown in Fig. 6. Traditional regression models, such as EDSR, RCAN, IRN, and HCFlow, tend to generate overly-smooth results with few sharp edges. Although GAN-based methods, including ESRGAN and VQIR, generate more details, their edge information lacks accuracy and regularity. In contrast, our approach, by leveraging the prior knowledge encapsulated in the Stable Diffusion model, is able to reconstruct more accurate and rich edge details.
4.3 Ablation Study
We conduct ablation experiments to validate the effectiveness of each component. All ablations are trained on the DF2K dataset and tested on the validation set of the DIV2K.
4.3.1 Effectiveness of Feature Rescaling
To verify the effectiveness of the feature rescaling operation, we move the rescaling operation from the latent space to the pixel space. This means we directly train a downsampling network to downscale the HR $x$ to LR $y$. Subsequently, $y$ is used as a condition to guide the generation process of the SD model. As shown the 1st row in Tab. 2, the rescaling operation in the latent space can improve reconstruction performance, as evidenced by lower LPIPS and DISTS.
In Fig. 7, we plot the LPIPS variation curves on the validation set for the two rescaling schemes. The pixel space-based rescaling scheme requires a longer time to converge, whereas the latent space based scheme can stably converge to a better result in a shorter time. Additionally, we observe that the pixel space based scheme experiences severe color shifts in the middle of training, as illustrated in Fig. 8. It requires a substantial number of iterations to learn the global color distribution of the reconstructed images. This is because using pixels as the conditioning input for Stable Diffusion makes it difficult to align with the intrinsic priors of the pre-trained model, leading to issues such as unstable training and longer training time.
Figure 7: The validation LPIPS curve for rescaling models based on latent space and pixel space.
Figure 8: Reconstructed results after 60k training steps. The pixel space based scheme shows severe color shifts.
Table 2: Ablation study of rescaling space and SD prior.
Latent space Pixel space SD prior LPIPS\ $\downarrow$ DISTS\ $\downarrow$
$\times$ $\checkmark$ $\checkmark$ 0.3818 0.1333
$\checkmark$ $\times$ $\times$ 0.4675 0.3109
$\checkmark$ $\times$ $\checkmark$ 0.3559 0.1188
4.3.2 Effectiveness of SD Prior
To validate that SD model provides rich priors about real-world HR images, we directly omit the second training stage. As shown in Tab. 2 (2nd row), the introduction of the SD priors can significantly improve performance of LPIPS and DISTS and greatly enhance the perceptual quality of the reconstruction results.
4.3.3 Effectiveness of DFRM Design
Our DFRM includes three specific designs: the domain converter module, the pixel guidance module, and the stop-gradient operation in the loss function Equ. 11. To analyze their impact, we remove each component individually and plot the validation loss curves, as shown in Fig. 10. It can be observed that the introduction of the INN and the stop-gradient operation can both reduce reconstruction error. Although the pixel guidance module does not improve reconstruction performance, our experiments show that it can enhance the fidelity and visual quality of the LR images, as shown in Tab. 3. Therefore, the design of our DFRM strikes a good balance between feature reconstruction and LR image quality.
Ideally, the domain converter composed of a set of invertible neural networks should be able to accurately perform bidirectional mapping between the feature domain and the pixel domain, enabling the LR to share the same content with HR images. However, due to the limited representation power of INN, the intermediate features $w$ obtained by simply downscaling the latent code $z$ are not well-suited for conversion to the pixel domain. The introduction of pixel guidance allows for embedding some pixel-domain information into the intermediate features $w$, thereby achieving a balance between reconstruction performance and LR image quality. As shown in Fig. 9, the LR images exhibit significant noise and color distortions in the absence of pixel guidance. However, the introduction of pixel guidance significantly alleviates these issues.
Figure 9: Comparisons of LR images visual quality with and without pixel guidance.
Table 3: Ablation study of pixel guidance module. The introduction of pixel guidance can significantly enhance the quality of LR images.
Method PSNR\ $\uparrow$ SSIM\ $\uparrow$ NIQE\ $\downarrow$
w/o pixel guidance 19.88 0.5898 22.33
with pixel guidance 21.99 0.7237 13.27
Figure 10: The validation loss curves when removing a specific component from the DFRM.
4.4 Discussions
4.4.1 Comparison with JPEG Compression
Our model can be combined with existing image compression methods to achieve more extreme compression ratios. To validate this, we use the lossless compression method PNG to compress LR images and compare its performance with JPEG compression at various LR sizes. Specifically, we train rescaling models at three different scales (8$\times$, 16$\times$, and 32$\times$) and record the bpp by the LR images. We also adjust the quality parameters of JPEG compression accordingly for comparison. As shown in Fig. 11, at the same bits per pixel (bpp), the image reconstruction quality of our model significantly surpasses that of JPEG compression. In Fig. 12, we present the visual quality, the bpp and the LPIPS of the reconstructed images. Compared to JPEG compression, our method requires comparable or less storage space while achieving superior objective metrics. Additionally, JPEG compression tends to produce noticeable block artifacts and color distortions in the background regions of the images, whereas our model maintains superior reconstruction quality across all three scales. These comparisons demonstrate the potential of our model for image compression.
Figure 11: R-D curves of JPEG compression and our proposed model on the DIV2K validation set.
Figure 12: Visual comparisons between our model and JPEG compression. Below the images, the bpp and LPIPS of the reconstructed images are provided.
4.4.2 Fidelity-realism Trade-off
Our two-stage model outputs two features in the latent space ($\hat{z}$ and $\hat{z}_{0}$), which represent reconstruction results with high fidelity and high realism respectively. It is important to note that the high fidelity mentioned here does not refer to higher pixel fidelity but to semantic fidelity in the latent space. By fusing the results of the two stages, we can achieve a trade-off between fidelity and realism:
$z_{\mathrm{fus.}}=\alpha\hat{z}+\left(1-\alpha\right)\hat{z}_{0}$ (17)
$x_{\mathrm{fus.}}=\mathcal{D}\left(z_{\mathrm{fus.}}\right)$ (18)
As shown in Fig. 13, by adjusting the value of $\alpha$, we can achieve a smooth transition between fidelity and realism. More results are provided in the supplementary materials. As shown in Fig. 14, It can be observed that by adjusting the fusion weights, we can achieve a smooth transition between fidelity and realism.
Figure 13: By adjusting the fusion ratio of two latent codes, we can achieve a trade-off between fidelity and realism.
Figure 14: By adjusting the fusion weights, our model can achieve a trade-off between fidelity and realism. The perceptually optimized results display richer details, but they also exhibit some artifacts.
Figure 15: Visual comparisons of LR images.
5 Limitations
Although our model demonstrates superior performance in extreme image rescaling tasks, it still has some limitations. One of them is the quality of the LR images. As shown in Fig. 15, the LR images generated by IRN closely resemble the bicubic downsampled image. In contrast, the LR images produced by HCFlow and VQIR exhibit some color shifts. LR images generated by our model are slightly noisy. In the future, more advanced invertible network architectures and more effective loss functions could be explored to further improve the quality of LR images.
6 Conclusion
Extreme image rescaling is a highly challenging yet significant task. In this paper, we propose a novel framework called LSBIR to address issues such as insufficient texture and semantic inaccuracies. Specifically, we leverage the rich priors encapsulated in the pre-trained SD model to mitigate the ill-posed nature of rescaling. In the first stage, we design a DFRM to achieve bidirectional mapping between HR features and LR images. In the second stage, the rescaled features from the first stage serve as conditions to guide the generation process of the SD model through an effective adapter. Both qualitative and quantitative experiments demonstrate the superiority of our approach.